# Controllable Video Generation: A Survey

### v1

@@ 6.3 Hybrid and Scalable Autoregressive Methods
Ensuring temporal coherence remains a key challenge in video generation. Diffusion models excel at producing high-quality individual frames but often struggle with smooth transitions and logical frame-to-frame progression, especially in long-duration videos. Autoregressive methods, which generate frames sequentially while conditioning on prior outputs, offer a solution for enforcing temporal dependencies. For example, NOVA [50] demonstrates the potential of these methods for generating temporally consistent videos. However, their high computational cost, particularly for long sequences or high resolutions, limits scalability.
Hybrid strategies that combine Autoregressive methods with diffusion models present a promising direction. Diffusion models could operate in latent spaces to generate consistent representations, which lightweight Autoregressive mechanisms refine to ensure temporal coherence. This approach balances efficiency with consistency, making it suitable for long-duration videos.
Another promising avenue is multi-scale modeling, where coarse temporal dynamics are captured at lower resolutions, and fine details are refined at higher resolutions. This technique, explored in works like Motion-I2V [196], is particularly effective for maintaining long-term consistency alongside high-quality local details.
To further enhance scalability, memory-efficient architectures such as sparse attention transformers or recurrent latent modules could be utilized to reduce the computational load of modeling long-range dependencies. Additionally, self-supervised pretraining on large video datasets like WebVid-10M [14] or Panda-70M [15] could improve generalization across tasks, minimizing the need for task-specific fine-tuning.
By integrating hybrid approaches and addressing computational challenges, future systems could generate videos that are both high-quality and temporally coherent, enabling applications in storytelling, simulation, and animation.

### v2

@@ 7.3 Hybrid and Scalable Autoregressive Methods
Scaling to long videos with strong temporal coherence remains difficult. Diffusion models produce high-quality frames but may suffer from accumulated inconsistencies over long durations, whereas autoregressive generation explicitly models temporal dependencies and can improve coherence (e.g., NOVA [deng2025autoregressivevideogenerationvector]), at the cost of higher latency and compute.
Hybrid designs that combine diffusion and autoregressive components are therefore attractive: diffusion can generate globally consistent latent representations, while lightweight autoregressive modules enforce sequential continuity. Multi-scale modeling is also promising, capturing coarse dynamics at low temporal/spatial resolution and refining details progressively (e.g., Motion-I2V [10.1145/3641519.3657497]). For further scalability, memory-efficient architectures (e.g., sparse attention, recurrent latent modules) can reduce long-range modeling cost, and large-scale self-supervised pretraining on datasets such as WebVid-10M [Bain21WebVid-10M] and Panda-70M [chen2024panda-70m] can improve generalization and reduce reliance on task-specific tuning.
Overall, advances in hybrid architectures and efficient long-range modeling are likely to be central to practical, high-quality, and controllable long-video generation for applications such as storytelling, simulation, and animation.
@@ 8 Conclusion
Data Availability Statements. All the data used in this paper can be downloaded from the Internet, and the authorization has been obtained, including civitai [civitai].